OpenStudy (anonymous):

what is the inverse of y=-3x+6/7

OpenStudy (anonymous):

interchange x and y then, find y

OpenStudy (mathmale):

I see ambiguity here. Do you mean y = (-3x+6)/7 or do you mean y=-3x + (6/7)? Please use parentheses for clarity.

OpenStudy (anonymous):

6 over 7 as in six sevenths

OpenStudy (anonymous):

y= (-3x+6)/7 now you switch x and y x= (-3y+6)/7 7x = -3y+6 7x-6+=-3y (7x-6)/-3 =y make sense?

OpenStudy (mathmale):

I'd suggest you start over then: Find the inverse of y=-3x + (6/7).

OpenStudy (anonymous):

f-1(x) = x/-3 + 18/7